# PETLP: A Privacy-by-Design Pipeline for Social Media Data in AI Research

### v1

@@ 5 Conclusion
This paper has introduced PETLP to address a fundamental challenge in contemporary AI research: harnessing social media data’s potential whilst respecting legal obligations and ethical responsibilities. Our framework repositions privacy and compliance from constraints to design principles that enhance research rigour and public trust. Through detailed legal analysis and practical implementation, we demonstrate that responsible research practice strengthens rather than compromises scientific innovation – creating conditions for sustainable, trustworthy AI development that withstands regulatory scrutiny and maintains social licence. We position PETLP not as a definitive solution but as a structured foundation for ongoing dialogue about how AI research can serve societal benefit whilst respecting individual privacy.
@@ Appendix A GDPR Principles for Social Media Research
This consideration is particularly significant when recognising that social media data frequently contains or enables the inference of ‘special categories of personal data’ under Article 9 GDPR. Such data includes information revealing or permitting the deduction of sensitive characteristics such as health status, racial or ethnic origin, religious beliefs, or political opinions. Recent CJEU rulings (CJEU 2022, 2023b, 2023a) have clarified that explicit disclosure of such characteristics is not required to trigger Article 9 protections. Instead, the mere possibility of inferring or deducing protected attributes – regardless of the accuracy of such inferences – activates these heightened legal safeguards. This principle carries profound implications for applied AI research, where advanced models may infer sensitive characteristics from seemingly innocuous patterns of language use, user interactions, or posting behaviour. AI profiling techniques can similarly transform ordinary personal data into special category data when combined with other information or contextual analysis (Article 29 Data Protection Working Party 2018).
@@
Formally, the research institution typically acts as the controller (European Data Protection Board 2020, p. 3). However, controllership is a functional concept determined by actual influence over data processing decisions, not merely formal designation (European Data Protection Board 2020). In practice, AI researchers decides data collection methodologies, transformation techniques, and model training approaches that directly impact data subjects’ privacy. These decisions – such as which data fields to collect, what anonymisation techniques to employ, or how to balance model performance against privacy preservation – are integral to determining the “means” of processing.
@@
Article 24 of the GDPR (“Responsibility of the Controller”) operationalises the accountability principle established in Article 5(2), requiring controllers to implement appropriate technical and organisational measures to ensure and demonstrate compliance. This accountability framework adopts a risk-based approach, with the DPIA serving as its primary mechanism when processing is ‘likely to result in a high risk to the rights and freedoms of natural persons’ (Article 35).
@@
Importantly, DPIAs are not static or one-off exercises. They are “living” documents that must be revisited and updated when there are changes to the processing activities, models, or data sources (Article 29 Data Protection Working Party 2017, p. 9) – an aspect particularly relevant in iterative AI development cycles where models are frequently re-trained or re-purposed. While the GDPR does not mandate full public disclosure of DPIAs, publishing a summary of the assessment and its key findings can serve as a powerful mechanism for fostering public trust and demonstrating organisational transparency.
@@
The availability of these bases depends fundamentally on the organisational context. Article 6(1)(e) permits processing when necessary for the performance of a task carried out in the public interest or in the exercise of official authority vested in the controller. Both public authorities and private entities can rely on this basis when performing legally-established public interest tasks. Universities, as public authorities whose constitutional documents typically define research as part of their role, generally rely on public interest for research activities. For example, University College London produced a Statement of Tasks in the Public Interest (UCL Data Protection Office 2018), ‘facilitating and carrying out research in any field’, which may include research entirely funded by private companies, stating that most of UCL’s research will be undertaken using the ‘public task’ condition. Notably, Article 6(1)(f) explicitly states that legitimate interests ‘shall not apply to processing carried out by public authorities in the performance of their tasks’. This prohibition means that whilst private companies and independent research organisations may invoke legitimate interests, public authorities conducting research within their official remit cannot. Researchers must distinguish clearly between their broader interest (the overarching objective, such as improving mental health detection) and their specific purpose (the concrete processing activity, such as analysing Reddit posts to train a classification model).
@@
Each legal basis imposes distinct procedural requirements. Public interest processing under Article 6(1)(e) requires demonstrating a clear basis in law for the task or authority, which need not be a specific statutory power but must be grounded in constitutional documents, charters, or legal frameworks. The processing must be necessary – meaning no less intrusive means exist to achieve the objective. When processing special category data, researchers must additionally satisfy Article 9(2)(j) by implementing Article 89(1) safeguards, including data minimisation, pseudonymisation where feasible, and ethical oversight. In contrast, legitimate interests under Article 6(1)(f) – available only to non-public authorities or public authorities acting outside their official tasks – demands a documented Legitimate Interest Assessment (LIA) that evaluates: (1) the legitimacy of the interest pursued; (2) the necessity of processing, confirming no less intrusive means exist; and (3) a balancing test weighing the controller’s interests against data subjects’ rights and freedoms. As emphasised in EDPB Guidelines 1/2024, only if this balancing test favours the controller may processing proceed.
@@
Regardless of the chosen legal basis, the GDPR affords scientific research privileged status through multiple flexibilities designed to facilitate socially beneficial research whilst maintaining privacy protections. Recital 159 mandates a broad interpretation of “scientific research”, encompassing ‘technological development and demonstration, fundamental research, applied research and privately funded research’, provided it follows recognised methodological and ethical standards (European Data Protection Board 2021). These privileges include purpose compatibility for secondary use (Article 5(1)(b)), extended retention periods (Article 5(1)(e)), transparency exemptions when individual notification requires disproportionate effort (Article 14(5)(b)), and conditional limitations on erasure and objection rights where these would seriously impair research objectives (Articles 17(3)(d) and 21(6)). However, these flexibilities do not exempt researchers from implementing appropriate Article 89(1) safeguards – rather, they represent a calibrated balance between enabling valuable research and protecting individual privacy through technical and organisational measures.
@@ Appendix B WP29 Criteria for High-Risk Processing Requiring DPIA
The GDPR does not provide a precise definition of “high risk” processing that mandates a DPIA. Instead, Article 35(3) offers non-exhaustive examples of processing operations likely to result in high risk:
@@ B.1 Application to Social Media Research
The WP29 guidelines establish that processing meeting two or more criteria typically requires a DPIA. Moreover, the more criteria are met by the processing, the more likely it is to present a high risk to the rights and freedoms of data subjects, and therefore to require a DPIA, regardless of the measures which the controller envisages to adopt (Article 29 Data Protection Working Party 2017, p. 11).
@@ Appendix C Integrating the EU Artificial Intelligence Act into PETLP
Although the most visible transparency duties of the EU Artificial Intelligence Act (“AI Act”) – such as content-origin disclosure and watermarking for generative systems (Article 50) – arise when a model or dataset is shared with external audiences, the Regulation embeds legally enforceable controls across all stages of the PETLP pipeline. This appendix examines how the AI Act’s requirements map onto our framework and addresses a critical tension between differential privacy techniques and data quality obligations.
@@ C.1 Mapping AI Act Requirements to PETLP Stages
Any AI system that will ultimately be placed on the market or put into service within the Union must, from its inception, operate under a risk-management framework (Article 9) and data-governance controls (Article 10). For providers who scrape or ingest Reddit, X, or other platform corpora, this means documenting collection processes, representativeness, potential biases and mitigation strategies before the first byte is written to disk. If the resulting model is likely to be classed as general-purpose AI (GPAI) (following Article 51), the provider must additionally prepare a “sufficiently detailed” summary of the training data categories (Annex XI, Article 53).
@@
During the Load phase, researchers configure storage, access controls and intra-consortium sharing. Article 12 obliges providers of high-risk systems to maintain automatic logs ’to ensure a level of traceability’ commensurate with the risk. When data travel outside the EEA, GDPR transfer rules (Articles 44–49) apply in parallel, and the AI Act’s ex-territorial reach ensures that non-EU collaborators who are “providers” or “deployers” must also respect these logging and documentation requirements.
@@
Article 2(6) excludes AI systems “developed and put into service for the sole purpose of scientific research and development” from the Act’s operational scope. Academic teams who never release their models, weights or APIs therefore remain governed solely by the GDPR-plus-IP bundle analysed earlier. Industrial R&D units, however, must re-enter the Act’s compliance track the moment a prototype is commercialised or otherwise “put into service”. Likewise, an open-source release escapes the GPAI chapter only if the model poses no “systemic risk” (Article 53(2)).
@@ C.2 Reconciling Differential Privacy with Article 10(3) Data-Quality Duties
Article 10(3) AI Act requires that training, validation and test datasets be “relevant, sufficiently representative and, to the best extent possible, free of errors and complete” in view of the system’s intended purpose. [^5] Official Journal, Reg. (EU) 2024/1689, Article 10(3). At first sight, the calibrated noise that constitutes differential privacy appears to inject errors into the data, thereby frustrating compliance. A closer reading shows the relationship to be one of managed trade-off rather than outright conflict:
@@
    Risk-based standard: ’Free of errors’ …to the best extent possible is a reasonableness clause, not a strict liability rule. Recital 67 that underpins Article 10(3) recognises that ensuring representativeness or completeness may require balancing competing constraints, including privacy protection.
@@
    Scope of the duty: Article 10(3) speaks to the statistical properties of the dataset as a whole. DP noise is, by design, zero-mean and unbiased; it degrades individual-level fidelity whilst preserving aggregate distributions within a quantifiable “privacy budget”. Hence a DP-protected corpus can remain “sufficiently representative” if the privacy parameter $\varepsilon$ is chosen so that downstream model performance meets documented acceptance criteria. The Act’s risk-management system (Article 9) provides the forum for justifying that calibration.
@@
    Complementarity with GDPR minimisation: Differential privacy operationalises GDPR data-minimisation by guaranteeing that no single record has material influence on the output; the AI Act does not override this principle but obliges providers to evidence that the resulting utility remains adequate.
@@ C.3 Additional Resources
These instruments complement the regulatory framework by translating abstract obligations into actionable guidance, particularly valuable for researchers navigating the intersection of AI Act requirements with existing GDPR and IP obligations. While not legally binding (except the template requirement under Article 53), they represent authoritative interpretations that supervisory authorities are likely to reference when assessing compliance.
@@ Appendix D Decision Trees for Legal Compliance in Social Media AI Research
This appendix provides ten decision trees operationalising the legal requirements discussed throughout this paper. Each addresses a specific compliance challenge in social media research. (Note: Trees contain detailed pathways optimised for digital viewing. Zoom in to see all decision nodes, citations, and guidance clearly.)

### v2

@@ 5 Conclusion
PETLP repositions privacy and compliance from constraints to design principles that enhance research rigour and public trust. Through detailed legal analysis and practical implementation, we demonstrate that responsible research practice strengthens rather than compromises scientific innovation – creating conditions for sustainable, trustworthy AI development that withstands regulatory scrutiny and maintains social licence. We position PETLP not as a definitive solution but as a structured foundation for ongoing dialogue about how AI research can serve societal benefit whilst respecting individual privacy in an era where these values increasingly conflict.
Acknowledgments
Dr. Giorgos D. Vrakas was supported by a postdoctoral fellowship within the ONISILOS MSCA COFUND program at the University of Cyprus, funded by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101034403.
Legal Disclaimer
This paper is intended for informational and academic purposes only. The statements made regarding the law, ethics, and logistics of social media data practices are based on general understanding and research as of the publication date. This document does not constitute legal advice, nor does it create an attorney-client relationship. Laws and regulations surrounding data extraction, processing, storage, and presentation may vary by jurisdiction and are subject to change. The authors make no representations or warranties regarding the accuracy, completeness, or applicability of the information contained herein. Readers are encouraged to consult qualified Data Protection Officers or legal professionals to obtain advice specific to their individual circumstances before implementing any of the approaches described. The authors disclaim all liability for any actions taken based on the content of this paper.
@@ Appendix A GDPR Principles for Social Media Research
This consideration is particularly significant when recognising that social media data frequently contains or enables the inference of ‘special categories of personal data’ under Article 9 GDPR. Such data includes information revealing or permitting the deduction of sensitive characteristics such as health status, racial or ethnic origin, religious beliefs, or political opinions. Recent CJEU rulings (Court of Justice of the European Union 2022, 2023b, 2023a) have clarified that explicit disclosure of such characteristics is not required to trigger Article 9 protections. Instead, the mere possibility of inferring or deducing protected attributes – regardless of the accuracy of such inferences – activates these heightened legal safeguards. This principle carries profound implications for applied AI research, where advanced models may infer sensitive characteristics from seemingly innocuous patterns of language use, user interactions, or posting behaviour. AI profiling techniques can similarly transform ordinary personal data into special category data when combined with other information or contextual analysis (Article 29 Data Protection Working Party 2018).
@@
Formally, the research institution typically acts as the controller (European Data Protection Board 2020, p. 3). However, controllership is a functional concept determined by actual influence over data processing decisions, not merely formal designation (European Data Protection Board 2020). In practice, AI researchers decides data collection methodologies, transformation techniques, and model training approaches that directly impact data subjects’ privacy. These decisions – such as which data fields to collect, what anonymisation techniques to employ, or how to balance model performance against privacy preservation – are integral to determining the means of processing.
@@
Article 24 of the GDPR (‘Responsibility of the Controller’) operationalises the accountability principle established in Article 5(2), requiring controllers to implement appropriate technical and organisational measures to ensure and demonstrate compliance. This accountability framework adopts a risk-based approach, with the DPIA serving as its primary mechanism when processing is ‘likely to result in a high risk to the rights and freedoms of natural persons’ (Article 35).
@@
Importantly, DPIAs are not static or one-off exercises. They are living documents that must be revisited and updated when there are changes to the processing activities, models, or data sources (Article 29 Data Protection Working Party 2017, p. 9) – an aspect particularly relevant in iterative AI development cycles where models are frequently re-trained or re-purposed. While the GDPR does not mandate full public disclosure of DPIAs, publishing a summary of the assessment and its key findings can serve as a powerful mechanism for fostering public trust and demonstrating organisational transparency.
@@
The availability of these bases depends fundamentally on the organisational context. Article 6(1)(e) permits processing when necessary for the performance of a task carried out in the public interest or in the exercise of official authority vested in the controller. Both public authorities and private entities can rely on this basis when performing legally-established public interest tasks. Universities, as public authorities whose constitutional documents typically define research as part of their role, generally rely on public interest for research activities. For example, University College London adopted a Statement of Tasks in the Public Interest (UCL Data Protection Office 2018) defining its mission as ‘facilitating and carrying out research in any field’. This broad mandate encompasses even research entirely funded by private companies, with UCL stating that most research activities fall under the ‘public task’ condition. Notably, Article 6(1)(f) explicitly states that legitimate interests ‘shall not apply to processing carried out by public authorities in the performance of their tasks’. This prohibition means that whilst private companies and independent research organisations may invoke legitimate interests, public authorities conducting research within their official remit cannot. Researchers must distinguish clearly between their broader interest (the overarching objective, such as improving mental health detection) and their specific purpose (the concrete processing activity, such as analysing Reddit posts to train a classification model).
@@
Each legal basis imposes distinct procedural requirements. Public interest processing under Article 6(1)(e) requires demonstrating a clear basis in law for the task or authority, which need not be a specific statutory power but must be grounded in constitutional documents, charters, or legal frameworks. The processing must be necessary – meaning no less intrusive means exist to achieve the objective. When processing special category data, researchers must additionally satisfy Article 9(2)(j) by implementing Article 89(1) safeguards, including data minimisation, pseudonymisation where feasible, and ethical oversight. In contrast, legitimate interests under Article 6(1)(f) – available to private entities and public authorities acting outside their official tasks – requires a documented Legitimate Interest Assessment (LIA) that evaluates: (1) the legitimacy of the interest pursued; (2) the necessity of processing, confirming no less intrusive means exist; and (3) a balancing test weighing the controller’s interests against data subjects’ rights and freedoms. Commercial research organisations, private consultancies, and industry research divisions typically rely on this basis, as they cannot invoke public task provisions. As emphasised in EDPB Guidelines 1/2024, only if this balancing test favours the controller may processing proceed.
@@
Regardless of the chosen legal basis, the GDPR affords scientific research privileged status through multiple flexibilities designed to facilitate research whilst maintaining privacy protections. Recital 159 mandates a broad interpretation of ‘scientific research’, encompassing ‘technological development and demonstration, fundamental research, applied research and privately funded research’, provided it follows recognised methodological and ethical standards (European Data Protection Board 2021). These privileges include purpose compatibility for secondary use (Article 5(1)(b)), extended retention periods (Article 5(1)(e)), transparency exemptions when individual notification requires disproportionate effort (Article 14(5)(b)), and conditional limitations on erasure and objection rights where these would seriously impair research objectives (Articles 17(3)(d) and 21(6)). However, these flexibilities do not exempt researchers from implementing appropriate Article 89(1) safeguards – rather, they represent a calibrated balance between enabling research and protecting individual privacy through technical and organisational measures.
@@ Appendix B WP29 Criteria for High-Risk Processing Requiring DPIA
The GDPR does not provide a precise definition of ‘high risk’ processing that mandates a DPIA. Instead, Article 35(3) offers non-exhaustive examples of processing operations likely to result in high risk:
@@ B.1 Application to Social Media Research
The WP29 guidelines establish that processing meeting two or more criteria typically requires a DPIA. Moreover, the more criteria are met by the processing, the more likely it is to present a high risk to the rights and freedoms of data subjects, and therefore to require a DPIA (Article 29 Data Protection Working Party 2017, p. 11). This assessment operates independently of any measures which the controller envisages to adopt.
@@ Appendix C Integrating the EU Artificial Intelligence Act into PETLP
Although the most visible transparency duties of the EU Artificial Intelligence Act (AI Act) – such as content-origin disclosure and watermarking for generative systems (Article 50) – arise when a model or dataset is shared with external audiences, the Regulation embeds legally enforceable controls across all stages of the PETLP pipeline. This appendix examines how the AI Act’s requirements map onto our framework and addresses a critical tension between differential privacy techniques and data quality obligations.
@@ C.1 Mapping AI Act Requirements to PETLP Stages
Any AI system that will ultimately be placed on the market or put into service within the Union must, from its inception, operate under a risk-management framework (Article 9) and data-governance controls (Article 10). For providers who scrape or ingest Reddit, X, or other platform corpora, this means documenting collection processes, representativeness, potential biases and mitigation strategies before the first byte is written to disk. If the resulting model is likely to be classed as general-purpose AI (GPAI) (following Article 51), the provider must additionally prepare a ‘sufficiently detailed’ summary of the training data categories (Annex XI, Article 53).
@@
During the Load phase, researchers configure storage, access controls and intra-consortium sharing. Article 12 obliges providers of high-risk systems to maintain automatic logs ‘to ensure a level of traceability’ commensurate with the risk. When data travel outside the EEA, GDPR transfer rules (Articles 44–49) apply in parallel, and the AI Act’s ex-territorial reach ensures that non-EU collaborators who are ‘providers’ or ‘deployers’ must also respect these logging and documentation requirements.
@@
Article 2(6) excludes AI systems ‘developed and put into service for the sole purpose of scientific research and development’ from the Act’s operational scope. Academic teams who never release their models, weights or APIs therefore remain governed solely by the GDPR-IP-contract bundle analysed earlier. Industrial R&D units, however, must re-enter the Act’s compliance track the moment a prototype is commercialised or otherwise ‘put into service’. Likewise, an open-source release escapes the GPAI chapter only if the model poses no ‘systemic risk’ (Article 53(2)).
@@ C.2 Reconciling Differential Privacy with Article 10(3) Data-Quality Duties
Article 10(3) AI Act requires that training, validation and test datasets be ‘relevant, sufficiently representative and, to the best extent possible, free of errors and complete’ in view of the system’s intended purpose. At first sight, the calibrated noise that constitutes differential privacy appears to inject errors into the data, thereby frustrating compliance. A closer reading shows the relationship to be one of managed trade-off rather than outright conflict:
@@
    Risk-based standard: ‘Free of errors’ …to the best extent possible is a reasonableness clause, not a strict liability rule. Recital 67 that underpins Article 10(3) recognises that ensuring representativeness or completeness may require balancing competing constraints, including privacy protection.
@@
    Scope of the duty: Article 10(3) speaks to the statistical properties of the dataset as a whole. DP noise is, by design, zero-mean and unbiased; it degrades individual-level fidelity whilst preserving aggregate distributions within a quantifiable privacy budget. Hence a DP-protected corpus can remain ‘sufficiently representative’ if the privacy parameter $\varepsilon$ is chosen so that downstream model performance meets documented acceptance criteria. The Act’s risk-management system (Article 9) provides the forum for justifying that calibration.
@@
    Complementarity with GDPR minimisation: Differential privacy operationalises GDPR data minimisation by guaranteeing that no single record has material influence on the output; the AI Act does not override this principle but obliges providers to evidence that the resulting utility remains adequate.
@@ C.3 Additional Resources
These instruments complement the regulatory framework, particularly valuable for researchers navigating the intersection of AI Act requirements with the PETLP framework. While not legally binding (except the template requirement under Article 53), they represent authoritative interpretations that supervisory authorities are likely to reference when assessing compliance.
@@ Appendix D Decision Trees for Legal Compliance in Social Media AI Research
This appendix provides nine decision trees operationalising the legal requirements discussed throughout this paper. Each addresses a specific compliance challenge in social media research.
Note: Trees contain detailed pathways optimised for digital viewing. Zoom in to see all decision nodes, citations, and guidance clearly.